OpenStudy (anonymous):

-4 times a number plus 16 is -166

OpenStudy (anonymous):

please write an equation

OpenStudy (anonymous):

@triciaal

OpenStudy (anonymous):

this is what i think -4 x n+16 but what does this mean 16 is -166

Nnesha (nnesha):

times/product = multiply quotient/divide = divide sum/increasing/morethan means add difference = subtract in the statement `is ` represent equal sign `=`

OpenStudy (anonymous):

so then it will be 4-xn+16 = -166

OpenStudy (anonymous):

sorry i mean -4xn+16=-166

Nnesha (nnesha):

yes here is an example \[2 \times n \] can be written as 2n

Directrix (directrix):

This part is okay. Just avoid using x as a multiplication symbol. -4 x n+16 Use -4n + 16 Add this part: is -166 Think of "is" as being this =

Nnesha (nnesha):

` -4xn` here multiplication sign looks like `x` variable so you can write it as -4n

OpenStudy (anonymous):

ohh ok Directrix and nnesha

OpenStudy (anonymous):

but how do i cheak if its wright using subtraction

Nnesha (nnesha):

wait it's a number +16

Directrix (directrix):

-4xn+16=-166 This is okay with one exception. We do not know if you are using x as a variable or multiplication symbol. -4n means -4 times the number n. You don't need the x to show multiplication. -4n+16 =-166 is all

OpenStudy (anonymous):

how do i cheack if it right using subtraction @Directrix @Nnesha

Directrix (directrix):

The instruction you gave is "write an equation." You did that. Now, are you saying that you also want to solve for n?

OpenStudy (anonymous):

yes sorry if i wasnt clear in the beginning my apologies

Directrix (directrix):

What do you get when you subtract 16 from both sides?